OpenStudy (anonymous):

write an equation of a line that passes through the points (-2,5) and (1,2) show work

OpenStudy (amistre64):

-2,5 1,2 ----- -3,3 = slope = -3/3 = -1 dint i just do this one like 5 minutes ago lol

OpenStudy (amistre64):

y = -x + 2

myininaya (myininaya):

yes you did

myininaya (myininaya):

you got y=-x+3

OpenStudy (amistre64):

slope is: change in y 3 --------- = --- = -1 change in x -3 y = -x + b 2 = -(1) + b 3 = b y = -x +3